OpenStudy (anonymous):

which part of the qu solved through factorizingadratic formula tells whether it can be

OpenStudy (solomonzelman):

You can factor the \(\large\color{blue}{ \rm ax^2+bx+c }\) if \(\large\color{blue}{ \rm b^2-4(a)(c)}\) (the discriminant) is a perfect square.

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

plz solve for x in 16x squared -81=0

OpenStudy (solomonzelman):

\(\large\color{blue}{ \rm 16x^2-81=0 }\) \(\large\color{blue}{ \rm 16x^2=81 }\) \(\large\color{blue}{ \rm 16x^2=9^2 }\) \(\large\color{blue}{ \rm 4^2x^2=9^2 }\) \(\large\color{blue}{ \rm (4x)^2=9^2 }\) \(\large\color{blue}{ \rm 4x=± 9 }\) \(\large\color{blue}{ \rm x=± 9/4 }\)

OpenStudy (anonymous):

i think you are a genius...thnx

OpenStudy (anonymous):

4x squared + 3 = 12x